OpenStudy (anonymous):

PLEASE HELP ME I'LL GIVE MEDALS What is the solution to the equation -2(1 + 5a) = -6(a+3) ?

OpenStudy (anonymous):

-2-10a=-6a-18 +2 =+2 -10a=-6a-16 +6a =+6a -4a =-16 /-4=/-4 a=4